# Search for the production of dark gauge bosons in the framework of Einstein-Cartan portal in the simulation of proton-proton collisions at $\sqrt{s}=13.6$ TeV

### v8

@@ V Results
The shape-based analysis has been used based on the missing transverse momentum distributions ($E^{\text{miss}}_{T}$), which are good discriminant variables since the signal distributions are characterized by relatively large $E^{\text{miss}}_{T}$ values compared to SM backgrounds. After applying selection 3 listed in table 3, the distribution of the missing transverse momentum is illustrated in figure 6 for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
Figure 6: The distribution of the missing transverse momentum, for events passing selection 3 presented in table 3, for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
@@
$N_{s}$ and $N_{b}$ represent the number of signals and the total number of SM background events that pass the selections (3) mentioned in tables 3. In figure 7, the significance of A discovery potential is plotted versus the integrated luminosity for the simplified model based on Einstein-Cartan gravity, with the dark matter mass fixed at $M_{\chi}=500$ GeV, $M_{ST}=2000$ GeV and the coupling constants $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$. In the run3-LHC experiment, there is a possibility of detecting evidence for the torsion field with an integrated luminosity of more than 9 fb-1 for $M_{ST}=2000$ GeV and $M_{A^{\prime}}=200$. Furthermore, for $M_{ST}=2000$ GeV and $M_{A^{\prime}}=700$, a $5\sigma$ discovery can be achieved with an integrated luminosity exceeding 38 fb-1, as illustrated in figure 7. However, according to the discovery potential based on the integrated luminosity, achieving a $5\sigma$ evidence is not feasible when the $M_{A^{\prime}}$ masses exceed 700 GeV due to the significant decrease in the signal cross sections.
@@
Figure 7: The significance was presented with the integrated luminosity for Einstein-Cartan gravity, considering different dark boson masses $(M_{A^{\prime}})$. The values of the coupling constants were fixed at $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$, with $M_{ST}=2000$ GeV, and the dark matter mass set at $M_{\chi}=500$ GeV. The red dotted line refers to the $5\sigma$ discovery potential.
@@
Figure 8: 95% CL upper limits on the cross-section times the branching ratio (expected), as a function of the mediator’s mass ($M_{ST}$) based on Einstein-Cartan model, with the muonic decay of the A. The black line represents the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV.
Figure 9: The 95% CL upper limits on the product of the cross-section and branching fraction from the inclusive search, for variations of pairs of the simplified model parameters ($M_{ST}$ and $M_{A^{\prime}}$). The filled region indicates the upper limit. The dotted black curve indicates the expected exclusions for the nominal A cross-section.
@@
To make a statistical interpretation for our results, we performed a statistical test based on the profile likelihood method, with the use of the modified frequentist construction CLs [58, 59] used in the asymptotic approximation [57] to derive exclusion limits on the product of signal cross sections and branching fraction Br($A^{\prime}$ $\rightarrow\mu\mu$) at 95% confidence level. The limits and significances presented have an adhoc flat 10% uncertainty systematic uncertainties taken into account to cover all possible systematic effects.
@@
The 95% upper limit on the cross-section times the branching ratio versus the mass of torsion field $M_{ST}$, for the simplified model based on Einstein-Cartan gravity, is presented in figure 8, with the muonic decay of the A and coupling constant values of $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$ and dark matter mass $M_{\chi}=500$ GeV. The black solid curve represents the simplified model for $M_{A^{\prime}}=200$ GeV. Based on figure 8, we exclude the torsion field $(ST)$ production in the mass range between 1369 - 6241 GeV as shown from the expected median.
@@
For the simplified model (based on Einstein-Cartan gravity), the cross-section times the branching ratio limit is presented in Figure 9 as a function of the mediator’s masses $M_{ST}$ and the masses of the dark neutral gauge boson $M_{A^{\prime}}$. The region between the respective pair of the expected 95% dotted line is excluded. The results from the inclusive signal regions exclude expected values of up to $62411369<M_{ST}<$ 6241 GeV.
@@ VI Summary
We have proposed a methodology to search for the dark neutral gauge bosons (A) at the LHC, which is produced in association with dark matter $(\chi)$ in the framework of $U(1)_{D}$ simplified model based on the Einstein-Cartan gravity. The presented analysis has been performed using the simulated proton-proton collisions corresponding to the LHC run 3 with 13.6 TeV center of mass energy, for an integrated luminosity of 52 fb-1, which corresponds to the amount of certified data collected by the CMS experiment so far during LHC run 3. Results from the muonic decay mode of A are discussed, with fixing the values of the coupling constants to be $g_{D}=1.2$, $g_{\eta}=0.2$, and dark matter mass ($M_{\chi}=500$ GeV). We have considered the variations of several parameters of the signal model: the torsion field mass $(M_{ST})$ and the dark neutral gauge boson mass $(M_{A^{\prime}})$.
The final analysis indicates that achieving a $5\sigma$ discovery of the A dark gauge boson in the $\mu^{+}\mu^{-}$ decay channel is feasible. This can be accomplished for $M_{A^{\prime}}$ values ranging from 200 to 700 GeV, given $M_{ST}=2000$ GeV, with an integrated luminosity of less than 40 fb-1 at the 13.6 TeV center of mass energy of the LHC run 3.
In case no new physics has been discovered, the general version of the analysis, which uses only event-level kinematic variables, excludes models with $62411369<M_{ST}<$ 6241 GeV at 95% confidence level (CL).
@@ Acknowledgements.
The author of this paper would like to thank Cao H. Nam, the author of [17], for his useful discussions about the theoretical models, and for sharing with us the Universal FeynRules Output (UFO) for the model that was used for the generation of the events. This paper is based on works supported by the Science, Technology, and Innovation Funding Authority (STIF) under grant number 48289.

### v9

@@ V Results
The shape-based analysis has been used based on the missing transverse energy distributions ($E^{\text{miss}}_{T}$), which are good discriminant variables since the signal distributions are characterized by relatively large $E^{\text{miss}}_{T}$ values compared to SM backgrounds. After applying selection 3 listed in table 3, the distribution of the missing transverse energy is illustrated in figure 6 for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
Figure 6: The distribution of the missing transverse energy, for events passing selection 3 presented in table 3, for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
@@
$N_{s}$ and $N_{b}$ represent the number of signals and the total number of SM background events that pass the selections (3) mentioned in tables 3.
@@
In figure 7, the significance of A discovery potential is plotted versus the integrated luminosity for the simplified model based on Einstein-Cartan gravity, with the dark matter mass fixed at $M_{\chi}=500$ GeV, $M_{TS}=2000$ GeV and the coupling constants $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$. In the run3-LHC experiment, there is a possibility of detecting evidence for the torsion field with an integrated luminosity of more than 9 fb-1 for $M_{TS}=2000$ GeV and $M_{A^{\prime}}=200$. Furthermore, for $M_{TS}=2000$ GeV and $M_{A^{\prime}}=700$, a $5\sigma$ discovery can be achieved with an integrated luminosity exceeding 38 fb-1, as illustrated in figure 7. However, according to the discovery potential based on the integrated luminosity, achieving a $5\sigma$ evidence is not feasible when the $M_{A^{\prime}}$ masses exceed 700 GeV due to the significant decrease in the signal cross sections.
@@
Figure 7: The significance was presented with the integrated luminosity for Einstein-Cartan gravity, considering different dark boson masses $(M_{A^{\prime}})$. The values of the coupling constants were fixed at $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$, with $M_{TS}=2000$ GeV, and the dark matter mass set at $M_{\chi}=500$ GeV. The red dotted line refers to the $5\sigma$ discovery potential.
@@
Figure 8: 95% CL upper limits on the cross-section times the branching ratio (expected), as a function of the mediator’s mass ($M_{\texttt{TS}}$) based on the Einstein-Cartan model for $\texttt{g}_{D}=1.0$ 8(a) and 1.2 8(b), with the muonic decay of the A. The solid colored curves illustrate the cross-section measurements multiplied by the branching ratios (in pb) for the Einstein-Cartan gravity simplified model with $\texttt{g}_{\eta}=$ 0.061 or 0.067 (in blue), 0.125 (in red), and 0.20 (in black).
@@
To make a statistical interpretation for our results, we performed a statistical test based on the profile likelihood method, with the use of the modified frequentist construction CLs [60, 61] used in the asymptotic approximation [59] to derive exclusion limits on the product of signal cross sections and branching fraction Br($A^{\prime}$ $\rightarrow\mu\mu$) at 95% confidence level. The limits and significances presented have an adhoc flat 10% uncertainty, systematic uncertainties taken into account to cover all possible systematic effects.
@@
The 95% upper limit on the product of the cross-section and the branching ratio, plotted against the mass of the torsion field $M_{TS}$, is shown in figure 8. This analysis is based on a simplified model using Einstein-Cartan gravity and incorporates the muonic decay of the $A^{\prime}$ for coupling constants $\texttt{g}_{D}=1.0$ 8(a) and $1.2$ 8(b). In this context, we consider a dark matter mass of $M_{\chi}=500$ GeV and $M_{A^{\prime}}=200$ GeV. The solid curves illustrate the cross-section measurements multiplied by the branching ratios (in pb) for a simplified model with $\texttt{g}_{\eta}=$ 0.061 or 0.067 (in blue), 0.125 (in red) and 0.20 (in black). These measurements correspond to various values of the torsion field mass ($M_{TS}$), as detailed in Table 2, and are calculated at a center-of-mass energy of $\sqrt{s}=13.6$ TeV.
In Figure 8, the intersections of the solid-colored curves with the median expected curve, represented as a dotted line, indicate the mass ranges where the production of the torsion scalar (TS) can be excluded. For a coupling constant of $\texttt{g}_{\eta}=0.125$, this exclusion range is between 1476 and 4860 GeV. For $\texttt{g}_{\eta}=0.2$, the range extends from 1387 to 6040 GeV when $\texttt{g}_{D}=1.0$. Similarly, when $\texttt{g}_{D}=1.2$, the ranges are as follows: from 1460 to 5100 GeV for $\texttt{g}_{\eta}=0.125$ and from 1369 to 6241 GeV for $\texttt{g}_{\eta}=0.2$.
Our findings show that $\texttt{g}_{\eta}$ must be greater than 0.067 for $\texttt{g}_{D}$ = 1 and 0.061 for $\texttt{g}_{D}$ = 1.2.
These findings are summarized in Table 4. As a result, when we set $M_{\chi}=500$ GeV and $M_{A^{\prime}}=200$ GeV, we find that larger values of the coupling constant $\texttt{g}_{\eta}$ (such as 0.2) lead to the exclusion of a wider range of torsion field mass values. On the other hand, when the $\texttt{g}_{D}$ coupling is increased from 1.0 to 1.2, the outcome remains relatively unchanged.
$\texttt{g}_{D}$ $\texttt{g}_{\eta}=0.125$ $\texttt{g}_{\eta}=0.2$
1.0 1476 - 4860 GeV 1387 - 6040 GeV
1.2 1460 - 5100 GeV 1369 - 6241 GeV
Table 4: The expected limit values at 95% CL for different values of the model coupling constants $\texttt{g}_{D}$ and $\texttt{g}_{\eta}$.
@@
In the context of the dynamic torsion field, the classical value of $\texttt{g}_{\eta}=1/8$ becomes unstable due to quantum corrections, leading to variations in $\texttt{g}_{\eta}$ with changes in the energy scale [17]. Consequently, we have opted for $\texttt{g}_{\eta}=0.2$. Thus, in our simplified model, which is based on Einstein-Cartan gravity and features coupling constants of $\texttt{g}_{D}=1.2$, $\texttt{g}_{\eta}=0.2$, and a dark matter mass of $M_{\chi}=500$ GeV, we present the cross-section times the branching ratio limit in Figure 9. This is illustrated as a function of the mediator masses $M_{TS}$ and the masses of the dark neutral gauge boson $M_{A^{\prime}}$. The region between the respective pair of the expected 95% dotted line is excluded. The results from the inclusive signal regions exclude expected values of up to $1369<M_{TS}<6241$ GeV.
Figure 9: The 95% CL upper limits on the product of the cross-section and branching fraction from the inclusive search, for variations of pairs of the simplified model parameters ($M_{\texttt{TS}}$ and $M_{A^{\prime}}$). The filled region indicates the upper limit. The dotted black curve indicates the expected exclusions for the nominal A cross-section.
@@ VI Summary
We have proposed a methodology for searching for the torsion field (TS) at the LHC, which can decay into dark matter (DM) pair particles. One of these DM particles is sufficiently heavy to decay into dark neutral gauge bosons (A) along with another DM particle. This study has been conducted within the framework of the $U(1)_{D}$ simplified model, based on Einstein-Cartan gravity.
The presented analysis has been performed using the simulated proton-proton collisions corresponding to the LHC run 3 with 13.6 TeV center of mass energy, for an integrated luminosity of 52 fb-1, which corresponds to the amount of certified data collected by the CMS experiment so far during LHC run 3.
Results from the muonic decay mode of A are discussed, with fixing the values of the dark matter mass ($M_{\chi}=500$ GeV). We have considered the variations of several parameters of the signal model: the torsion field mass $(M_{TS})$ and the dark neutral gauge boson mass $(M_{A^{\prime}})$ in addition to the model coupling constants $\texttt{g}_{\eta}$ and $\texttt{g}_{D}$.
The final analysis indicates that achieving a $5\sigma$ discovery of the A dark gauge boson in the $\mu^{+}\mu^{-}$ decay channel is feasible. This can be accomplished for $M_{A^{\prime}}$ values ranging from 200 to 700 GeV, given $M_{TS}=2000$ GeV, with an integrated luminosity of less than 40 fb-1 at the 13.6 TeV center of mass energy of the LHC run 3.
In our current analysis, we conducted limit calculations at a 95% confidence level for both the cross section and the torsion mass ($M_{TS}$), specifically with values set at $M_{\chi}=500$ GeV and $M_{A^{\prime}}=200$ GeV. We not only established limits on the cross section and torsion mass but also examined a variety of values for the model coupling constants, $\texttt{g}_{\eta}$ and $\texttt{g}_{D}$.
Our results show that $\texttt{g}_{\eta}$ needs to be greater than 0.067 for $\texttt{g}_{D}$ = 1 and 0.061 for $\texttt{g}_{D}$ = 1.2; notably, higher values of this coupling constant, such as 0.2, lead to the exclusion of a wider range of torsion field mass values. Furthermore, our findings suggest that adjusting $\texttt{g}_{D}$ to either 1.0 or 1.2 results in only a minimal effect on the limit calculations.
If no new physics emerges, the comprehensive analysis, which is based solely on event-level kinematic variables, excludes models within the mass range of $1369<M_{TS}<6241$ GeV at a 95% confidence level. This applies to $M_{A^{\prime}}$ values spanning from 200 to 900 GeV, given that $\texttt{g}_{\eta}$ is set to 0.2 and $\texttt{g}_{D}$ is at 1.2.
@@ Acknowledgements.
The author of this paper would like to thank Cao H. Nam, the author of [17], for his useful discussions about the theoretical models, and for sharing with us the Universal FeynRules Output (UFO) for the model that was used for the generation of the events. This paper is based on works supported by the Science, Technology, and Innovation Funding Authority (STDF) under grant number 48289.